You invested 14,000 in two accounts paying 6% and 9% annual interest, respectively. If the total interest earned for the year $900, how much was invested at each rate?
Answer by rfer(16322)

0.06x+0.09(14000-x)=900
0.06x+1260-0.09x=900
-0.03x=-360
x=$12,000 at 6%
14000-12000=$2,000 at 9%